OpenStudy (anonymous):

Maria placed a beaker in her yard to measure the amount of rain that fell during the school week. She checked the beaker everyday and found that 1/4 of an inch fell everyday for 3 days. On the 4th day there was no change. On the fifth day the amount of rain in the beaker had decreased by 1/8 of an inch. How much rain in inches was in Maria's beaker on the 5th day? Express your answer in decimal form.

OpenStudy (anonymous):

1/4+1/4+1/4+0-1/8

OpenStudy (anonymous):

so 1/4+ 1/4 + 1/4 +1/4 * 1/4 is 5/4

OpenStudy (anonymous):

1/4 for 3 days = 1/4 x 3 = 3/4 No change on the 4th day so it is still 3/4 decreases by 1/8 on 5th day = 3/4 - 1/8 3/4 - 1/8 (find common denominator which is 8) 6/8 - 1/8 = 5/8 5/8 in decimal form is .625 ANSWER : .625 inches were in Maria's beaker on the fifth day :)